OpenStudy (anonymous):

Solve the equation 8D-4D-6D-8=2D A.) -1 B.) 0 C.) -4 D.) -2

OpenStudy (anonymous):

8d-4d-6d=2d+8 -2d -2d 6d-4d-6d=8 2d-6d= -4d -4d=8 8/-4= -2 d= -2